OpenStudy (anonymous):

(1+cos alpha)/(sin alpha) +(sin alpha)/(1+cos alpha) Simplify

OpenStudy (anonymous):

\[\frac{ 1+\cos \alpha }{ \sin \alpha }+\frac{ \sin \alpha }{ 1+\cos \alpha}\]

OpenStudy (anonymous):

I got it down to \[\frac{ 2 }{ \sin \alpha(1+\cos \alpha) }\]

OpenStudy (anonymous):

Not sure thats right but >.<

OpenStudy (whpalmer4):

what if you try multiplying top and bottom by the conjugate of \((1+\cos \alpha)\)?

OpenStudy (anonymous):

k so that would end up like 2/sin alpha

OpenStudy (anonymous):

which would be 2csc x

OpenStudy (anonymous):

which is an answer.

OpenStudy (anonymous):

and holy crap its right too! :O You're magic.

OpenStudy (anonymous):

OKAY problem solved, onto the next one. I'll try to figure it out before i post it again.

OpenStudy (whpalmer4):

multiplying by the conjugate is another trick worth remembering...

OpenStudy (whpalmer4):

it's the (x-y)(x+y) = (x^2-y^2) bit all over again

OpenStudy (anonymous):

okay. will try to remember/ understand.

OpenStudy (whpalmer4):

It pops up in lots of places: Simplifying radicals \[\frac{1}{1+\sqrt{2}} = \frac{1}{1+\sqrt{2}}*\frac{1-\sqrt{2}}{1-\sqrt{2}}=\frac{1-\sqrt{2}}{1^2-(\sqrt{2})^2} = \frac{1-\sqrt{2}}{-1} = \sqrt{2}-1\] Complex numbers \[\frac{1}{a+bi} = \frac{1}{a+bi}*\frac{a-bi}{a-bi} =\frac{a-bi}{a^2-b^2i^2} =\frac{a-bi}{a^2+b^2}\] and so on

OpenStudy (anonymous):

I may have to send you a gift basket after tonight my friend.

OpenStudy (whpalmer4):

If you get the hang of doing this, that will be reward enough...but next time you have some pizza, eat an extra slice for me!

OpenStudy (anonymous):

:) I just ate pretty much a whole homemade calzone in stress.